OpenStudy (anonymous):

Two women leave a shopping center by car, traveling in opposite directions. If one car averages 45 mph and the other 50 mph, how long will it take for the cars to be 285 apart?

OpenStudy (anonymous):

285 miles? Then 3 hours, as they move apart at a speed of 95 mi/h

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

can you answer another question for me, please?

OpenStudy (anonymous):

Jessica takes four tests in an english class. on each successive test, her score improves by 4 points. If her average score was 66%, what did she score on her first test?

OpenStudy (anonymous):

she scored 60

OpenStudy (anonymous):

can you explain that to me please?

OpenStudy (anonymous):

Ok, let x be the score she got in the first test :) now, on the second, she got x + 4, in the third x + 8 and the fourth x + 12, right? So the average of that is: (x + (x+4) + (x+8) + (x+12)) / 4 = 66 And (x + (x+4) + (x+8) + (x+12)) / 4 = (4x + 24)/4 = x + 6 = 66, so x = 60